OpenStudy (anonymous):

Sarah wants to hang a mirror in her room but the mirror and frame must not have an area larger than 15 square feet. The mirror is 3 feet wide and 4 feet long. Which quadratic equation represents the area of the mirror and frame combined? (The frame must have equal width of x on each side.)

OpenStudy (anonymous):

where are the quadratic equations?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

aren't they meant to give u some? it says "Which quadratic equation represents ....." so it gives the indication that there should be some options like multiple choice.

OpenStudy (anonymous):



OpenStudy (anonymous):

3x2 + 14x − 2 = 0 4x2 + 14x − 3 = 0 x2 + 12x − 14 = 0 2x2 + 7x − 14 = 0

OpenStudy (anonymous):